OpenStudy (anonymous):

Verify the identity. tan (x + (pi/2)) = -cot x

OpenStudy (anonymous):

\[\tan \left( x +\frac{ \Pi }{ 2 } \right)=-\cot x\]

OpenStudy (anonymous):

break tan into sin and cos

OpenStudy (anonymous):

how do I do that

OpenStudy (anonymous):

tanx = (sin x)/ (cos x)

OpenStudy (anonymous):

hold on I will give you the solution

OpenStudy (anonymous):

thanks, but will you show the work so I can try the next one myself?

OpenStudy (whpalmer4):

If you use the definition of \(\tan x = \sin x/\cos x\) you can rewrite the problem as \[\tan(x+\frac{\pi}{2}) = \frac{\sin(x+\frac{\pi}{2})}{\cos(x+\frac{\pi}{2})}\] Next, use the formulas for the sin and cos of a sum of angles: \[\sin(A+B) = \sin A\cos B + \cos A\sin B\]\[\cos(A+B) = \cos A\cos B - \sin A\sin B\] to simplify that expression. Finally, remember that \[\cot x = \frac{1}{\tan x} = \frac{\cos x}{\sin x}\]

hartnn (hartnn):

or you could use tan (A+B) formula \(\tan (A+B)= \dfrac{\tan A +\tan B}{1-\tan A tan B}\) plug in A =x, B= pi/2

OpenStudy (jdoe0001):

or use the cofunction identities of $$ sin\pmatrix{x+\frac{\pi}{2}} = cos(x)\\ cos\pmatrix{x+\frac{\pi}{2}} = -sin(x) $$

OpenStudy (anonymous):

you woulf find the solution in the attachment

OpenStudy (anonymous):

@hartnn the problem with that formula is that tan (pi/2) is infinity

hartnn (hartnn):

yes, yes! thats why my next step was to divide the numerator and denominator by 'tan B' but that would be \(\large if\) the asker would have tried it, got stuck and asked...

OpenStudy (anonymous):

ok you make sense I just did it in my mind nice one

OpenStudy (anonymous):

@hartnn